OpenStudy (anonymous):

A parabola contains the points (0,0) , (-1,-2) and (1,6). What is the equation of parabola in standard form?

OpenStudy (anonymous):

work a matrix: for ax^2 + bx + c = y

OpenStudy (anonymous):

a0 + b0 + c = 0 ; c = 0 a1 - b = -2 a1 + b = 6 so it really reduces in this case to a system of 2 equations .... which is boring enough